# Multilevel neural simulation-based inference

### v2

@@ A.1 Proof of Theorem˜1
$\left\|\nabla_{z}f_{\phi}^{l}(z)-\nabla_{z}f_{\phi}^{l-1}(z)\right\|_{2}$
$=\left\|\nabla_{z}\log q^{\text{NLE}}_{\phi}\left(g^{l}(z)\right)-\nabla_{z}\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\right\|_{2}$
$=\Big{\|}\nabla_{z}g^{l}(z)\nabla\log q^{\text{NLE}}_{\phi}(g^{l}(z))-\nabla_{z}g^{l-1}(z)\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\Big{\|}_{2}$ (26)
$\leq\Big{\|}\nabla_{z}g^{l}(z)\left(\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l}(z)\right)-\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\right)\Big{\|}_{2}$
$\quad+\Big{\|}\left(\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\right)\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\Big{\|}_{2}$ (27)
$\leq\Big{\|}\nabla_{z}g^{l}(z)\Big{\|}_{2}\Big{\|}\nabla\log q^{\text{NLE}}_{\phi}(g^{l}(z))-\nabla\log q^{\text{NLE}}_{\phi}(g^{l-1}(z))\Big{\|}_{2}$
$\quad+\Big{\|}\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\Big{\|}_{2}\Big{\|}\nabla\log q^{\text{NLE}}_{\phi}(g^{l-1}(z))\Big{\|}_{2}$ (28)
Here, ˜26 follows from the chain rule, ˜27 follows by adding and subtracting the term $\nabla_{z}g^{l}(z)\nabla\log q^{\text{NLE}}_{\phi}(g^{l-1}(z))$ and using the triangle inequality, and ˜28 follows from the Cauchy-Schwarz inequality of the 2-norm. Squaring both sides of this inequality and taking expectations, we then obtain:
$\mathbb{E}\left[\left\|\nabla_{z}f_{\phi}^{l}(z)-\nabla_{z}f_{\phi}^{l-1}(z)\right\|^{2}_{2}\right]$
$\leq\mathbb{E}\Big{[}\Big{(}\Big{\|}\nabla_{z}g^{l}(z)\Big{\|}_{2}\Big{\|}\nabla\log q^{\text{NLE}}_{\phi}(g^{l}(z))-\nabla\log q^{\text{NLE}}_{\phi}(g^{l-1}(z))\Big{\|}_{2}$
$\quad+\Big{\|}\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\Big{\|}_{2}\Big{\|}\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\Big{\|}_{2}\Big{)}^{2}\Big{]}$
$\leq\mathbb{E}\Big{[}2\Big{\|}\nabla_{z}g^{l}(z)\Big{\|}^{2}_{2}\Big{\|}\nabla\log q^{\text{NLE}}_{\phi}(g^{l}(z))-\nabla\log q^{\text{NLE}}_{\phi}(g^{l-1}(z))\Big{\|}_{2}^{2}$
$\qquad+2\Big{\|}\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\Big{\|}_{2}^{2}\Big{\|}\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\Big{\|}_{2}^{2}\Big{]}.$ (29)
$\leq 2\mathbb{E}\left[\Big{\|}\nabla_{z}g^{l}(z)\Big{\|}^{4}_{2}\right]^{\frac{1}{2}}\mathbb{E}\left[\Big{\|}\nabla\log q^{\text{NLE}}_{\phi}(g^{l}(z))-\nabla\log q^{\text{NLE}}_{\phi}(g^{l-1}(z))\Big{\|}_{2}^{4}\right]^{\frac{1}{2}}$
$\qquad+2\mathbb{E}\left[\Big{\|}\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\Big{\|}_{2}^{4}\right]^{\frac{1}{2}}\mathbb{E}\left[\Big{\|}\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\Big{\|}_{2}^{4}\right]^{\frac{1}{2}}.$ (30)
where ˜29 follows from that fact that for $a,b\in\mathbb{R}$, we have $(a+b)^{2}\leq 2a^{2}+2b^{2}$, and ˜30 follows from the Cauchy-Schwarz inequality for expectations. To conclude this proof, we notice that the derivation from ˜14 to ˜20 can be modified by replacing $G^{0}$ by $G^{l}$ to get:
@@
where the last inequality holds thanks to Assumption (A2). Similarly, replacing $G^{0}$ by $G^{l}-G^{l-1}$ gives
$\mathbb{E}\left[\left\|\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\right\|^{4}_{2}\right]\leq K_{\text{grad}}\left(\left\|G^{l}-G^{l-1}\right\|^{4}_{W^{1,4}}+1\right).$ (32)
@@
$\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\right]$ $\leq K_{\text{score}}(\phi)\left(\left\|G^{l-1}\right\|_{W^{1,4}}^{4}+1\right)$ (33)
@@
$\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l}(z)\right)-\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\right]$ (35)
$=\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l}(z)\right)-\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\Big{|}\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}\geq\delta\right]$
$\qquad\quad\times\mathbb{P}\left[\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}\geq\delta\right]$
$\quad+\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l}(z)\right)-\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\Big{|}\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}<\delta\right]$
$\qquad\quad\times\mathbb{P}\left[\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}<\delta\right]$ (36)
$\leq\frac{16}{\delta}\left(\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l}(z)\right)\right\|^{4}_{2}\right]+\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\right]\right)\mathbb{E}\left[\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}\right]$
@@
$\leq K_{\text{score-diff}}(\phi,\delta)\Big{\|}G^{l}-G^{l-1}\Big{\|}_{W^{1,4}}^{4}$ (38)
where ˜36 follows due to the law of total expectation. For ˜37, the bound on the first term follows due to the triangle inequality and Markov’s inequality, and the bound on the second term follows due to the local-Lipschitz condition (i.e. Assumption (A3)) and the fact that a probability is always upper bounded by $1$. Finally, ˜38 follows using ˜34 and the definition of Sobolev norm, and by grouping all constants. Combining all of the above (i.e. ˜7, ˜9, ˜28, ˜31, ˜32, ˜34, and ˜38), we end up with
@@
$\qquad\qquad+K_{\text{grad}}^{\frac{1}{2}}\left(\left\|G^{l}-G^{l-1}\right\|^{4}_{W^{1,4}}+1\right)^{\frac{1}{2}}K_{\text{score}}(\phi)^{\frac{1}{2}}\left(S_{l-1}^{4}+1\right)^{\frac{1}{2}}\Big{)}$
$\leq\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}+1\right),$
where $K_{l}(\phi)$ combines all constants. This proves our second result and therefore concludes our proof.
Note that the proof follows analogously for NPE with the alternative locally-Lipschitz and growth condition assumptions. ∎
@@ A.2 Proof of Theorem˜2
$\text{Var}\left[{\ell}_{\text{MLMC}}(\phi);n_{0},\ldots,n_{L}\right]\leq\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|^{4}_{W^{1,4}}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}+1\right),$
where once again we write $\|\cdot\|_{W^{1,4}}$ to denote the weighted norm $\|\cdot\|_{W^{1,4}(\pi\times\mathbb{U})}$. Overall, we would like to solve the following optimisation problem:
@@
$:=\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|^{4}_{W^{1,4}}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}+1\right)$
@@
$=\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|^{4}_{W^{1,4}}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}+1\right)$
@@
$\frac{\partial\mathcal{L}(n_{0},\ldots,n_{L},\nu)}{\partial n_{l}}=-K_{l}(\phi)\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}+1\right)n_{l}^{-2}+\nu(C_{l}+C_{l-1})=0$
$\quad\Leftrightarrow\quad n^{\star}_{l}=\sqrt{\frac{K_{l}(\phi)}{\nu(C_{l}+C_{l-1})}\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}.$ (41)
@@
$\frac{\partial L(n_{0},\ldots,n_{L},\nu)}{\partial\nu}=n_{0}C_{0}+\sum_{l=1}^{L}n_{l}(C_{l}+C_{l-1})-C_{\text{budget}}=0$
@@
Plugging the results of ˜40 and ˜41 into ˜42, we get:
@@
$\quad+\sum_{l=1}^{L}\sqrt{\frac{K_{l}(\phi)}{\nu(C_{l}+C_{l-1})}\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}\times(C_{l}+C_{l-1})$
@@
$\quad\left.+\sum_{l=1}^{L}\sqrt{K_{l}(\phi)(C_{l}+C_{l-1})\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}\right),$
@@
$\nu=\frac{\left(\sqrt{K_{0}(\phi)C_{0}\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)}+\sum_{l=1}^{L}\sqrt{K_{l}(\phi)(C_{l}+C_{l-1})\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}+1\right)}\right)^{2}}{C_{\text{budget}}^{2}}.$ (43)
We can now use the expression for $\nu$ that we obtained in ˜43 to obtain a simplified expression for $n_{0},\ldots,n_{L}$ (using ˜40 and ˜41):
$n_{0}^{\star}$ $\propto\frac{C_{\text{budget}}}{\sqrt{C_{0}}}\sqrt{\left\|G^{0}\right\|_{W^{1,4}}^{4}+1},\quad\quad n_{l}^{\star}$ $\propto\frac{C_{\text{budget}}}{\sqrt{C_{l}+C_{l-1}}}\sqrt{\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1}\quad\text{for }l\in\{1,\ldots,L\}.$

### v4

@@ A.1 Proof of Theorem˜1
$\left\|\nabla_{z}f_{\phi}^{l}(z)-\nabla_{z}f_{\phi}^{l-1}(z)\right\|_{2}$ $=\left\|\nabla_{z}-\log\tilde{q}_{\phi}\left(g^{l}(z)\right)-\left(\nabla_{z}-\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\right)\right\|_{2}$
$=\left\|\nabla_{z}\log\tilde{q}_{\phi}\left(g^{l}(z)\right)-\nabla_{z}\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\right\|_{2}$
$=\Big\|\nabla_{z}g^{l}(z)\nabla\log\tilde{q}_{\phi}(g^{l}(z))-\nabla_{z}g^{l-1}(z)\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\Big\|_{2}$ (26)
$\leq\Big\|\nabla_{z}g^{l}(z)\left(\nabla\log\tilde{q}_{\phi}\left(g^{l}(z)\right)-\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\right)\Big\|_{2}$
$\quad+\Big\|\left(\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\right)\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\Big\|_{2}$ (27)
$\leq\Big\|\nabla_{z}g^{l}(z)\Big\|_{2}\Big\|\nabla\log\tilde{q}_{\phi}(g^{l}(z))-\nabla\log\tilde{q}_{\phi}(g^{l-1}(z))\Big\|_{2}$
$\quad+\Big\|\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\Big\|_{2}\Big\|\nabla\log\tilde{q}_{\phi}(g^{l-1}(z))\Big\|_{2}$ (28)
Here, ˜26 follows from the chain rule, ˜27 follows by adding and subtracting the term $\nabla_{z}g^{l}(z)\nabla\log\tilde{q}_{\phi}(g^{l-1}(z))$ and using the triangle inequality, and ˜28 follows from the Cauchy-Schwarz inequality of the 2-norm. Squaring both sides of this inequality and taking expectations, we then obtain:
$\mathbb{E}\left[\left\|\nabla_{z}f_{\phi}^{l}(z)-\nabla_{z}f_{\phi}^{l-1}(z)\right\|^{2}_{2}\right]$ $\leq\mathbb{E}\Big[\Big(\Big\|\nabla_{z}g^{l}(z)\Big\|_{2}\Big\|\nabla\log\tilde{q}_{\phi}(g^{l}(z))-\nabla\log\tilde{q}_{\phi}(g^{l-1}(z))\Big\|_{2}$
$\quad+\Big\|\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\Big\|_{2}\Big\|\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\Big\|_{2}\Big)^{2}\Big]$
$\leq\mathbb{E}\Big[2\Big\|\nabla_{z}g^{l}(z)\Big\|^{2}_{2}\Big\|\nabla\log\tilde{q}_{\phi}(g^{l}(z))-\nabla\log\tilde{q}_{\phi}(g^{l-1}(z))\Big\|_{2}^{2}$
$\quad+2\Big\|\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\Big\|_{2}^{2}\Big\|\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\Big\|_{2}^{2}\Big].$ (29)
$\leq 2\mathbb{E}\left[\Big\|\nabla_{z}g^{l}(z)\Big\|^{4}_{2}\right]^{\frac{1}{2}}\mathbb{E}\left[\Big\|\nabla\log\tilde{q}_{\phi}(g^{l}(z))-\nabla\log\tilde{q}_{\phi}(g^{l-1}(z))\Big\|_{2}^{4}\right]^{\frac{1}{2}}$
$\qquad+2\mathbb{E}\left[\Big\|\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\Big\|_{2}^{4}\right]^{\frac{1}{2}}\mathbb{E}\left[\Big\|\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\Big\|_{2}^{4}\right]^{\frac{1}{2}}.$ (30)
where ˜29 follows from that fact that for $a,b\in\mathbb{R}$, we have $(a+b)^{2}\leq 2a^{2}+2b^{2}$, and ˜30 follows from the Cauchy-Schwarz inequality for expectations. To conclude this proof, we notice that the derivation from ˜13 to ˜16 can be modified by replacing $G^{0}$ by $G^{l}$ gives:
@@
where the last inequality holds thanks to Assumption (A2). Similarly, replacing $G^{0}$ by $G^{l}-G^{l-1}$ and following the derivations from ˜13 to ˜20 gives
$\mathbb{E}\left[\left\|\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\right\|^{4}_{2}\right]\leq K_{\text{grad}}\left\|G^{l}-G^{l-1}\right\|^{4}_{W^{1,4}}.$ (32)
We notice that we lose the additive term since the last columns of the matrix $\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)$ form a zero matrix. This is because
$\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)$ (33)
$=\begin{bmatrix}\nabla_{u}G_{\theta}^{l}(u_{1})-\nabla_{u}G_{\theta}^{l-1}(u_{1})\cdots\nabla_{u}G_{\theta}^{l}(u_{m})-\nabla_{u}G_{\theta}^{l-1}(u_{m})&\nabla_{u}\theta-\nabla_{u}\theta\\ \nabla_{\theta}G_{\theta}^{l}(u_{1})-\nabla_{\theta}G_{\theta}^{l-1}(u_{1})\cdots\nabla_{\theta}G_{\theta}^{l}(u_{m})-\nabla_{\theta}G_{\theta}^{l-1}(u_{m})&\nabla_{\theta}\theta-\nabla_{\theta}\theta\end{bmatrix}$
$=\begin{bmatrix}\nabla_{u}G_{\theta}^{l}(u_{1})-\nabla_{u}G_{\theta}^{l-1}(u_{1})\cdots\nabla_{u}G_{\theta}^{l}(u_{m})-\nabla_{u}G_{\theta}^{l-1}(u_{m})&\textbf{0}\\ \nabla_{\theta}G^{l}_{\theta}(u_{1})-\nabla_{\theta}G^{l-1}_{\theta}(u_{1})\cdots\nabla_{\theta}G^{l}_{\theta}(u_{m})-\nabla_{\theta}G^{l-1}_{\theta}(u_{m})&\textbf{0}\end{bmatrix}$
The non-zero lower-right block, i.e., $\nabla_{\theta}\theta$ in ˜14, was the cause of the additive term in ˜20 and the final result (the upper-right block is always zero since $\nabla_{u}\theta=\textbf{0}$), which is now cancelled out.
@@
$\mathbb{E}\left[\left\|\nabla\log q_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\right]$ $\leq K_{\text{score}}(\phi)\left(\left\|G^{l-1}\right\|_{W^{1,4}}^{4}+1\right)$ (34)
@@
$\mathbb{E}\left[\left\|\nabla\log\tilde{q}_{\phi}\left(g^{l}(z)\right)-\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\right]$ (36)
$=\mathbb{E}\left[\left\|\nabla\log\tilde{q}_{\phi}\left(g^{l}(z)\right)-\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\Big|\left\|g^{l}(z)-g^{l-1}(z)\right\|_{2}^{4}\geq\delta\right]$
$\qquad\quad\times\mathbb{P}\left[\left\|g^{l}(z)-g^{l-1}(z)\right\|_{2}^{4}\geq\delta\right]$
$\quad+\mathbb{E}\left[\left\|\nabla\log\tilde{q}_{\phi}\left(g^{l}(z)\right)-\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\Big|\left\|g^{l}(z)-g^{l-1}(z)\right\|^{4}_{2}<\delta\right]$
$\qquad\quad\times\mathbb{P}\left[\left\|g^{l}(z)-g^{l-1}(z)\right\|_{2}^{4}<\delta\right]$ (37)
$\leq\frac{8}{\delta}\left(\mathbb{E}\left[\left\|\nabla\log\tilde{q}_{\phi}\left(g^{l}(z)\right)\right\|^{4}_{2}\right]+\mathbb{E}\left[\left\|\nabla\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\right\|^{4}_{2}\right]\right)\mathbb{E}\left[\left\|g^{l}(z)-g^{l-1}(z)\right\|^{4}_{2}\right]$
@@
$\leq\left(\frac{8}{\delta}K_{\text{score}}(\phi)(S_{l-1}^{4}+S_{l}^{4}+2)+K_{\text{Lip}}(\phi)^{4}\right)\mathbb{E}\left[\left\|g^{l}(z)-g^{l-1}(z)\right\|^{4}_{2}\right]$ (39)
$\leq K_{\text{score-diff}}(\phi,\delta)\Big\|G^{l}-G^{l-1}\Big\|_{W^{1,4}}^{4}.$ (40)
Here, ˜37 follows due to the law of total expectation. For ˜38, the bound on the first term follows due to $\|a-b\|_{2}^{4}\leq 8(\|a\|_{2}^{4}+\|b\|_{2}^{4})$ and Markov’s inequality, and the bound on the second term follows due to the local-Lipschitz condition (i.e. Assumption (A3)) and the fact that a probability is always upper bounded by $1$. Then, ˜39 follows using ˜35. Finally, ˜40 follows by grouping all constants together and noting that
$\mathbb{E}\left[\left\|g^{l}(z)-g^{l-1}(z)\right\|^{4}_{2}\right]$ $=\mathbb{E}\left[\left(\sum_{i=1}^{m}\|G^{l}(u_{i})-G^{l-1}(u_{i})\|^{2}_{2}\right)^{2}\right]$ (41)
$\leq m\sum_{i=1}^{m}\mathbb{E}\left[\|G^{l}(u_{i})-G^{l-1}(u_{i})\|^{4}_{2}\right]\leq m\|G^{l}-G^{l-1}\|_{W^{1,4}}^{4}$ (42)
Combining all of the above (i.e. ˜7, ˜9, ˜28, ˜31, ˜32, ˜35, and ˜39), we end up with
@@
$\qquad\qquad+K_{\text{grad}}^{\frac{1}{2}}\left(\left\|G^{l}-G^{l-1}\right\|^{4}_{W^{1,4}}\right)^{\frac{1}{2}}K_{\text{score}}(\phi)^{\frac{1}{2}}\left(S_{l-1}^{4}+1\right)^{\frac{1}{2}}\Big)$
$\leq\frac{K_{l}(\phi)}{n_{l}}\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}},$
where $K_{l}(\phi)$ combines all constants. This proves our second result and therefore concludes our proof. ∎
@@ A.2 Proof of Theorem˜2
$\text{Var}\left[{\ell}_{\text{MLMC}}(\phi);n_{0},\ldots,n_{L}\right]\leq\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|^{4}_{W^{1,4}}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}},$
where once again we write $\|\cdot\|_{W^{1,4}}$ to denote the norm $\|\cdot\|_{W^{1,4}(\pi\times\mathbb{U})}$. Overall, we would like to solve the following optimisation problem:
@@
$:=\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|^{4}_{W^{1,4}}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}$
@@
$=\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|^{4}_{W^{1,4}}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}$
@@
$\frac{\partial\mathcal{L}(n_{0},\ldots,n_{L},\nu)}{\partial n_{l}}=-K_{l}(\phi)\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}n_{l}^{-2}+\nu(C_{l}+C_{l-1})=0$
$\quad\Leftrightarrow\quad n^{\star}_{l}=\sqrt{\frac{K_{l}(\phi)}{\nu(C_{l}+C_{l-1})}\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}}.$ (45)
@@
$\frac{\partial\mathcal{L}(n_{0},\ldots,n_{L},\nu)}{\partial\nu}=n_{0}C_{0}+\sum_{l=1}^{L}n_{l}(C_{l}+C_{l-1})-C_{\text{budget}}=0$
@@
Plugging the results of ˜44 and ˜45 into ˜46, we get:
@@
$\quad+\sum_{l=1}^{L}\sqrt{\frac{K_{l}(\phi)}{\nu(C_{l}+C_{l-1})}\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}}\times(C_{l}+C_{l-1})$
@@
$\quad\left.+\sum_{l=1}^{L}\sqrt{K_{l}(\phi)(C_{l}+C_{l-1})\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}}\right),$
@@
$\nu=\frac{\left(\sqrt{K_{0}(\phi)C_{0}\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)}+\sum_{l=1}^{L}\sqrt{K_{l}(\phi)(C_{l}+C_{l-1})\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}}\right)^{2}}{C_{\text{budget}}^{2}}.$ (47)
We can now use the expression for $\nu$ that we obtained in ˜47 to obtain a simplified expression for $n_{0},\ldots,n_{L}$ (using ˜44 and ˜45):
$n_{0}^{\star}$ $\propto\frac{C_{\text{budget}}}{\sqrt{C_{0}}}\sqrt{\left\|G^{0}\right\|_{W^{1,4}}^{4}+1},\quad\quad n_{l}^{\star}$ $\propto\frac{C_{\text{budget}}}{\sqrt{C_{l}+C_{l-1}}}\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}\quad\text{for }l\in\{1,\ldots,L\}.$